If a sample of water absorbs 41.0 kJ of heat and the temperature of the sample changes from 25.0∘C to 56.0∘C, what is the mass of the sample in grams?

Remember that the specific heat capacity of water is 4.184Jg⋅∘C.
Answer #1

q    = 41Kj   = 41000J

t1   = 250C

t2   = 560C

c   = 4.184J/g-0C

q   = mct

41000   = m*4.184*(56-25)

m         = 41000/(4.184*31)   = 316.1g

mass of water = 316.1g >>>>answer